This week saw the return of the LRS2 but in its full glory, ie. both the B and R components. The instrument cooled down quickly and we were ahead of schedule by the third day. We did have a few problems with calibrations lamps, communications with the TCS, and inexperienced operators but all of that is expected when you get FIRST LIGHT with a science instrument on an effectively brand new telescope.
One of the highlights of the observing was moving back and forth between LRS2-B and LRS2-R, an observing mode we hoped to achieve eventually but were able to take advantage of in this first official commissioning run. The Austin team will be back in the beginning of March. In the mean time the LRS2 instruments remain at the telescope pumped down and cold to test the stability of the cryogenic system.
Happy new year. We did conduct some engineering operations over the holidays. Most of this involved small changes to the pointing and collecting data to remove the remaining guider drift we continue to see. We made some changes to the lowest order of the pointing corrections.
We have also been able to move specific stars into the guide probes and wave front sensors. This past week was the first time we actually had stars in two guide probes and in the wavefront sensor at the same time.
At the beginning of January we began commissioning of the calibration wavefront sensor using the deployable wavefront sensor in the central IFU position. This is done with geo-stationary satellites and our procedure is to align the primary mirror and then setup on the deployable wavefront sensor, null out any tip, tilt and focus remaining and then image through the calibration wavefront sensor. Once we have done that a few times back and forth we offset the telescope so that we can image the satellite with the operation wavefront sensor 576″ away. We can now do this repeatably.
The day staff has been getting electronics setup in the 2nd VIRUS enclosure and dealing with small problems that have been popping up with the tracker and other sub-systems. The most critical of these was an uncontrolled warm up of the LRS-r. This was caused by a fault in the liquid nitrogen system and the normal warnings that normally go out were shunted into a dead phone line. The LRS-2 was sent back to Austin for diagnostics and new safeties will be put in place. We hope to have LRS-2 back out here next week.
We have hired a new Telescope Operator and he is being trained. We do have a open position for a computer system admin person currently posted so if you have anyone who wants to get away from the hustle and bustle of city life but still work with cutting edge equipment have them look at the UT job posting.